OpenStudy (anonymous):

Help with homework please :(///a student makes random guesses on 6 multiple-choice questions. One of the 5 choices in each question is correct. What is the probability that the student will get at most 4 answers right?

OpenStudy (anonymous):

okay this question is quite abstract but here goes what i think A b c d e f A b c d e f A b c d e f A b c d e f A b c d e f A b c d e f imagine all the As are correct to get all correct, you'd need a probability of (1/5)^6, am i right?

OpenStudy (anonymous):

Hi there! I got the part p=1/5 (.2); n=6. What throws me off is the AT MOST 4 is correct.

OpenStudy (anonymous):

I am using a graphing calculator: binompdf (6,0.2,x); I need to find x. it is not 4 since the question didn't state EXACTLY 4, but AT MOST 4

OpenStudy (anonymous):

yeah i encountered that query too haha well, i'd suggest finding out P(getting at most 2 wrong), then subtract that probability from 1, otherwise it'd be very tedious

ganeshie8 (ganeshie8):

P(X <= 4) = 1 - P(X>4) = 1 - [P(X=5) + P(X=6)]

OpenStudy (dan815):

6 questions,5choices each p(0)+p(1)+p(2)+p(3)+p(4) = at most 4 right or 1-p(5)-p(6)

OpenStudy (dan815):

p(6) is an easy one there is only 1 way to get everything right 1 out of 5^6

OpenStudy (dan815):

p(5) 

OpenStudy (dan815):

and just 1 way to pick the 5 other choices

OpenStudy (dan815):

6*(1/5^5)

OpenStudy (dan815):

1-(p(5)+p(6) 1-(6/5^5 + 1/5^6)

OpenStudy (dan815):

here another way to do with probabililty

OpenStudy (dan815):

binomial thm

OpenStudy (dan815):

prob of getting right = 1/5 prob of getting 5 success out of 6 is then (6Choose5)(1/5)^5*(4/5)

OpenStudy (dan815):

ucan see this also works right away for 6 6Choose6*1/5^6 * 1 = 1/5^6

OpenStudy (dan815):

1-((6Choose5)(1/5)^5*(4/5) + 6Choose6*1/5^6 * 1)

OpenStudy (dan815):

i think i did something wrong while calculaing the prob(5) because its not quite matching up with the other method

ganeshie8 (ganeshie8):

\[ P(X\le 4) = 1 - \left[\binom{6}{5} (1/5)^5(4/5)^1 + \binom{6}{6} (1/5)^6(4/5)^0\right]\]